OpenStudy (anonymous):

An equilateral triangle has perimeter 66. What is the length of its altitude? A. 22 B. 22 square root 3 C. 11 square root 3 D. 11

OpenStudy (abb0t):

equilateral means all same sides. and if it's perimeter (total) has 66. Then, one side must also be \(\sf \frac{66}{3}\)

OpenStudy (anonymous):

so it'd be A?

OpenStudy (anonymous):

its very simple.. square root of (33*33)-(16.5*16.5)

OpenStudy (anonymous):

The answer comes lyk 28.57

OpenStudy (anonymous):

but thats wrong.. considering its not even in the awnser choices

OpenStudy (anonymous):

The altitude makes a right angle with one of the sides and half the base. Use that and the fact that a side is 22, along with c^2=a^2 + b^2.

OpenStudy (anonymous):

dear its not wrong u just need 2 do it in other form.... i'm new here so i'm not able to write it in an equation form... the basic is u need 2 first find the measure of one side and then use the pythagoras theorm

OpenStudy (anonymous):

first do what abb0t says n then follow pythagoras..... draw a diagram u'll get it.

OpenStudy (anonymous):

so its A

OpenStudy (anonymous):

nope dear... from the that you get 22 rite? then u do this 22^2=11^2+height^2 solve this n u'll get it.. sorry my fist answer was wrong!

OpenStudy (anonymous):

i still dont get it, but thanks anyways

OpenStudy (anonymous):

k.... good luck.. can you give me some help in biology?

OpenStudy (anonymous):

how do i see your biology question

OpenStudy (anonymous):

just send me a messege

OpenStudy (anonymous):

i'll ask it here wait... how does variation occur in asexual reproduction?

OpenStudy (anonymous):

someones typing on ur page